OpenStudy (anonymous):

What is the equation in point-slope form of the line passing through (0, 2) and (1, 5)?

OpenStudy (cgreenwade2000):

point slope form is \[y-y _{1}=m(x-x _{1})\]

OpenStudy (cgreenwade2000):

So, you are missing slope correct? How do you find slope using two points?

OpenStudy (anonymous):

Probably graph it and rise/run?

OpenStudy (cgreenwade2000):

You can graph it.... OR use slope formula: \[m= (y _{2}- y _{1})/(x _{2}-x _{1})\]

OpenStudy (anonymous):

so 0-1 and 2-5? so 1,3? or 1,-3?

OpenStudy (anonymous):

y – 5 = 3(x – 1) y – 2 = –3(x – 0) y + 2 = 3(x + 1) y + 5 = –3(x + 1)

OpenStudy (anonymous):

those are my choices though

OpenStudy (cgreenwade2000):

(0-1)/(2-5) 0-1=-1 2-5=?